Suppose that the sum of two whole numbers is 20, and the sum of their reciprocals is 5/16. Find the numbers.
Answer by stanbon(75887)

Suppose that the sum of two whole numbers is 20, and the sum of their reciprocals is 5/16. Find the numbers.
----------------------
Equations:
x + y = 20
1/x + 1/y = 5/16
-----------
x + y = 20
(x+y)/xy = 5/16
-----
20/xy = 5/16
5xy = 320
xy = 64
--
Remember x + y = 20
x = 20-y
---
Substitute:
(20-y)y = 64
20y-y^2 = 64
---------------
y^2-20y+64 = 0
y^2-16y-4y+64 = 0
y(y-16)-4(y-16) = 0
(y-16)(y-4) = 0
---
If y = 4; x = 16
If y = 16; x = 4
=======================
Cheers,
Stan H.